if k=(u-a)/s
and k(s) = u-a
then what do I do to the formula above to get a alone?
a = u - k(s) this is the answer but I don't know what to do to "u" to get it on the other side of the equation.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

if k=(u-a)/s
and k(s) = u-a (I don't see the relevance of this)
then what do I do to the formula above to get a alone?
---------------
k=(u-a)/s
ks = u-a
-a = ks-u
a = u - ks
------------


Answer by stanbon(75887)

if k=(u-a)/s
and k(s) = u-a
Solve for "a".
-------------------
Solve each equation for "u":
u-a = ks
u-a = k(s)
-----
u = a+ks
u = a+k(s)
------
It appears the following is true:
a = u-ks
a = u-k(s)
=================
ks must equal k(s)
Cheers,
Stan H.
=================